First system power on.

Fastlink connection seems broken between switch bcu and crate 0 . Mario reported a possible cable damage during the mounting.

=========================

Power:

set overcurrent for zenith:

print, set_curr_limit(60,0,/PUSHTOTAL)

Bypass chiuso, Cabinet Chiuso

P=3.3 bar

F=11.0 l/min

Cabinet 1 giro chiuso

P=3.2bar

F=10.9 l/min 10.3 l/min

Cabinet chiuso

P=3.3bar

F=11.0 l/min

Cabinet aperto 1 giro

P=3.2bar

F=10.3 l/min

Left to the above configuration for the cooling.

circa 03:31:40 cambiate le running mean di Flusso (100s->0s) e le umidita' (100s->10s)

Test di set_flat all'orizzonte

ok: oversampling just 50Hz (?)

the dome has been moved around and telescope have been slewed from horizon to zenith. No reset.

after the end of telescope movements, we set the oversampling to 500Hz

07:57:43: da circa 1 minuto fa oscillazioni sulla temperatura del cooling

fine dei test: Mon Mar 8 08:06:09 GMT 2010

-- MarcoXompero - 07 Mar 2010
Topic revision: r3 - 08 Mar 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback